# Coordinated Disclosure for AI: Beyond Security Vulnerabilities

### v2

@@ Abstract
Harm reporting in the field of Artificial Intelligence (AI) currently operates on an ad hoc basis, lacking a structured process for disclosing or addressing algorithmic flaws. In contrast, the Coordinated Vulnerability Disclosure (CVD) ethos and ecosystem play a pivotal role in software security and transparency. Globally, there are ongoing efforts to establish frameworks that promote transparency and collaboration in addressing AI-related issues, though challenges persist. Algorithmic flaws in machine learning (ML) models present distinct challenges compared to traditional software vulnerabilities, warranting a specialized approach. To address this gap, we propose the implementation of a dedicated Coordinated Flaw Disclosure (CFD) framework tailored to the intricacies of machine learning and artificial intelligence issues. This paper delves into the historical landscape of disclosures in ML, encompassing the ad hoc reporting of harms and the emergence of participatory auditing. By juxtaposing these practices with the well-established disclosure norms in cybersecurity, we argue that the broader adoption of CFD has the potential to enhance public trust through transparent processes that carefully balance the interests of both organizations and the community.
@@ 1 Introduction
On December 27, 2023, The New York Times filed a lawsuit against OpenAI and Microsoft for copyright infringement, alleging that ChatGPT was trained on the Times’ published works without authorization, providing evidence that ChatGPT generated articles verbatim (Mac 2023). This action contributed to a growing series of lawsuits against AI providers (O’Brien 2023; Authors Guild 2023; Fenwick & West LLP 2023; ACLU 2022). The specific instances referenced in The New York Times’ lawsuit are challenging to replicate, and as non-parties to the case, we lack insight into whether this is attributable to selective presentation by The New York Times or a true systematic memorization issue in ChatGPT’s design. The Times lawsuit has garnered significant public attention, at least partly due to The New York Times’ financial resources and extensive media reach – factors often unavailable to smaller harmed parties. This highlights the influence of such factors in shaping public discourse, which, in turn, has the potential to improve AI systems via outside pressure (Raji and Buolamwini 2022). Unexpected model behaviors, outside of its intended scope and purpose, underscore the need for robust frameworks to identify and address these issues effectively. This stands in contrast to the Common Vulnerabilities and Exposures (CVE) ecosystem, where vulnerabilities in software systems are systematically cataloged and addressed, often with greater clarity and consensus within the cybersecurity community.
@@
Historically, AI accountability initiatives have relied on periodic audits, emphasizing repetitive assessments but lacking a structured reporting framework for user-identified issues post-deployment, with recent policy efforts acknowledging this issue (European Commission 2020). This audit-centric paradigm is reflected in influential policies such as the U.S. Executive Order on AI (The White House 2023a), the EU’s draft AI Act (European Commission 2021), and New York City’s Local Law 144 (New York City Department of Consumer and Worker Protection 2021). However, this approach falls short when compared to the more comprehensive Coordinated Vulnerability Disclosure (CVD) processes standard in software security.
@@
Coordinated Vulnerability Disclosure (CVD) plays a crucial role as a mechanism for independent researchers to report newly identified vulnerabilities to affected vendors and the public (MultiMedia LLC 1999). This process enables transparent remediation before potential exploitation by malicious actors and has become a vital practice enshrined in government regulations and industry standards. Notably, the FDA mandates the implementation of CVD programs for medical device companies to enhance cybersecurity (U.S. Food and Drug Administration 2023).
@@
While CVD has demonstrated effectiveness in traditional software security, its direct application to machine learning (ML) systems faces unique challenges. ML diverges from traditional software in two key aspects: 1) ML model issues must satisfy statistical validity thresholds, and 2) concerns related to trustworthiness and bias extend beyond the typical scope of security vulnerabilities (Office of Science and Technology Policy 2022). It is important to note that numerous critical bias reports discussed on social media and in the news may not align with the accepted definition of a vulnerability [^1] The Coordinated Vulnerabilities Enumeration program, managed by MITRE and central to the CVD ecosystem, defines a vulnerability as: A problem in software, firmware, hardware, or service component resulting from a weakness that can be exploited, causing a negative impact on the confidentiality, integrity, or availability of the affected component or components..
@@
Upon recognizing this limitation in the conventionally accepted definition of a vulnerability, we propose to expand the terminology with a new proposed term - Flaw. We define a flaw as any unexpected model behavior that is outside of the defined intent and scope of the model design.
@@
In response to the unique challenges posed by machine learning (ML) systems, we advocate for the adaptation of the Coordinated Vulnerability Disclosure (CVD) process into a dedicated ”Coordinated Flaw Disclosure” (CFD) framework. Tailored to ML’s distinctive properties, CFD aims to formalize the recognition of valid issues in ML models through an adjudication process. This process seeks to strike a balance between the interests of vendors and the broader community, ultimately enhancing public trust by curating disclosed flaws into actionable reports.
@@
In the subsequent sections of the paper, we delve into specific aspects of this proposal. Section 2 provides essential background information, drawing a comparison between periodic audits and ad hoc harm reporting, while also exploring emerging participatory auditing practices. Section 3 examines evolving trends in disclosure practices within ML systems and identifies their shortcomings. Section 4 outlines the proposed changes within the CFD framework to improve disclosure rigor. Section 5 explores the practical implementation of CFD, proposing the utilization of model cards to define scope and intent. The paper concludes by emphasizing the potential of CFD to fortify AI accountability.
@@ 2 Background
In the following, we describe the background of our work. We will use the terminology introduced in the glossary in Table 1.
@@ 2.1 Coordinated Disclosure Conventions in Cybersecurity
CVD operates within a broader framework of programs that identify, index, and communicate security vulnerabilities, including the National Vulnerability Database, CVE system, and Common Vulnerability Scoring System (CVSS). These programs support diverse security operations such as vulnerability management, risk assessments, penetration testing, patch management, and threat intelligence (Cybersecurity Coalition 2019). CVE Numbering Authorities (CNA) are organizations external to MITRE that are authorized to issue CVEs within their scope, but MITRE acts as a ”root” that handles disputes. For example vulnerabilities in open source projects managed by the Apache Software Foundation will be sent to them first and then, if needed, appealed to MITRE (Apache 2024). In the EU, the European Cybersecurity Scheme on Common Criteria (EUCC) is currently developed to cover information and communications technology (ICT) based on the international Common Criteria scheme (European Union Agency for Cybersecurity (2024) ENISA).
@@ 3 Disclosure Practices in Machine Learning Systems: Evolving Trends
In this section, we examine the evolving trends of disclosure practices for machine learning systems and work that remains to be done for harmonization of these conventions. We break our analysis down into two axes – objective security and privacy flaw disclosures, and subjective ethical flaw disclosures, such as societal bias.
@@ 3.1 Security Disclosures for ML Systems
ML-based systems can be subject to three main types of disclosures from users and researchers: security vulnerability reports, model design failures, and error reports. In the following discussion, we examine examples under each category. Among these reports, some highlight actionable issues warranting recognition by external authorities, while others delve into details that may not demand significant attention.
@@
These reports highlight security vulnerabilities in ML models that are independent of the model’s performance. These are not usually model vulnerability reports, but are in components around the model. They often stem from traditional vulnerabilities in the components, in which case they should be referred to the Common Vulnerabilities and Exposures Numbering Authority (CNA) responsible for that specific component.
CVE reports on Numpy (CVE 2019a), Tensorflow (CVE 2019b), and Pytorch (CVE 2023c) fall into this category. The OpenAI bug bounty on Bugcrowd also addresses infrastructure flaws or vulnerabilities, as issues in the AI itself are considered out of scope (OpenAI 2023). Numerous examples exist of vulnerabilities that model inference pipelines might face, as the machine learning community often prioritizes reproducibility of experiments over code security. Efforts to address this concern are emerging, such as SafeTensors by Hugging Face (Hugging Face 2022).
An instance of a traditional security vulnerability caused by an ML model is DeepSloth (Hong et al. 2020), presenting a Denial Of Service attack similar to SlowLoris (Cloudflare 2024). This type of attack could be executed by an external entity on a Large Language Model (LLM) using early exit or multi-exit strategies to minimize costs. Consequently, this may lead to overprovisioning, reaching capacity, or system failure. Several model optimization techniques used by LLMs in production could be potential targets for such attacks.
A novel vulnerability is ProofPudding (MITRE 2019), the first CVE issued for an issue caused by a machine learning model. Model stealing, though concerning, demands numerous queries, especially for large models with complex behaviors. The risk is significantly reduced when the API provides excessive information. Proofpoint returned the exact output to attackers, minimizing the number of queries needed. However, they failed to implement rate limits on the information. The mitigation applied by ProofPoint was to limit the information their API provides, a practice theoretically possible with explainable AI techniques like Shapley values (Oksuz, Halimi, and Ayday 2023). The implementation of rate limits and limited information in responses represents best practices in ML security, and we anticipate more reports of this nature in the future.
@@
Model design failures are typically identified by manipulating samples to induce changes in the output. While they may resemble security vulnerabilities, unlike infrastructure reports, these vulnerabilities rely on modifying the output of a deployed machine learning model. They are not addressable simply improving the underlying dataset (Hooker 2021) and are shared among all models of a similar type. The complexity of these issues may render them unsolvable with current technology, leading to reports of this kind being considered out of scope or low priority.
The first documented vulnerability of this type in a production system was a “Bayesian Poisoning Attack” reported in 2003 on a bulletin board about PoPFile (The MIT 2004 Spam Conference ), and later on Paul Graham’s Blog (Graham 2003). This easily reproducible bypass did not affect the content of the spam but was valuable due to its focus on the method. Another model exhibiting similar behavior was the Cylance malware detector (Cyber 2019). Additional examples relevant to security practitioners come from the Malware Detection Evasion Competition, designed to elicit reports of these types (Anderson 2019). The competition rules included banning known bypasses like packing and awarded prizes for the most novel and interesting techniques (Anderson and Balazs 2020).
One frequently discussed model design failure is Adversarial Examples, which many machine learning security researchers consider low priority at best (Szegedy et al. 2014). These examples are often delicate, requiring a high level of technical expertise and control over the underlying system. Adversarial examples frequently do not survive jpeg compression (Aydemir, Temizel, and Temizel 2018) or a simple blur & repair (Carlini et al. 2022). Although one cited use for this attack is in self-driving cars, such scenarios usually demand extensive compromise of the car’s vision system. Adversarial patches represent a potential attack path (Brown et al. 2018) but are challenging to execute (Athalye et al. 2018). Marines tasked with bypassing an ML model monitoring CCTV footage succeeded with a low-complexity attack by getting into a large box (Scharre 2023b). In most scenarios, lower-hanging fruit for attackers is more appealing (Apruzzese et al. 2022), often resulting in these issues being considered out of scope or low priority for most ML systems, as vendors prioritize other pressing concerns.
@@
These cases involve issues that could stem from the expected statistical error rate demonstrated by the model during testing or systematic errors that are rectifiable through changes to the training data or modifications in how the model is trained. Such reports are ubiquitous, and vendors often incorporate telemetry and remediation tools into their systems for data gathering.
  • One-off Errors: These errors are an anticipated aspect of the model’s functioning. ML models seldom achieve 100% accuracy, and these errors are within the normal error rate and comply with the Service Level Agreement (SLA) when considered in conjunction with the rest of the query data.
  • Systematic Errors: These errors arise from unintended statistical biases in the model, leading to failure in producing output aligned with the intent of the model in one or more categories.
One-off errors are commonplace, such as spam ending up in the main inbox or legitimate emails being marked as spam. Actions like moving a legitimate email out of spam or reporting spam contribute to telemetry, enhancing subsequent model datasets. While these errors are typically perceived as expected inconveniences in using machine learning systems, they become reportable if they are egregious and cause harm. An extreme case is Google’s image classification system mislabeling a black person as a gorilla, which should be publicly reported using a coordinated disclosure process even if it occurs once (Scharre 2023a).
Systematic errors pose greater challenges in diagnosis, often originating from statistical biases in the dataset. These biases may result from socioeconomic forces, poor collection techniques, or unforeseen reasons during the model production, leading to harm. A classic example often shared with new machine learning students involves a tank classifier learning to recognize tanks based on the sky. British tanks were photographed with gray skies, and Russian tanks with blue skies, causing the classifier to learn the easier task of sky classification. While this specific story may be an urban legend, it serves as a cautionary tale for students (Branwen 2023). Another recent example of a systematic error is the “divergence attack” (Nasr et al. 2023) demonstrated by Google researchers on ChatGPT – when ChatGPT was asked to repeat a random word like “poem” infinitely, it started leaking private training data.
@@ 3.2 Disclosure of Ethical Issues in ML Systems
Despite calls for audits from academics, lawmakers and activists alike (Raji et al. 2022; Fjeld et al. 2020; Wilkinson et al. 2023), the identification and disclosure of ethical issues in machine learning (ML) systems have, for the past decade, progressed in a somewhat ad hoc and reactive manner. Unlike traditional software systems, which follow established practices for coordinated vulnerability disclosure, the process of surfacing problems with ML models has been inconsistent at best. Due to the lack of a mandated reporting mechanism like the one we outline in this paper, harms caused by flawed ML systems are often only discovered after the damage is done (Scharre 2023a). Few proactive, systematic mechanisms exist for auditing these complex sociotechnical systems, identifying risks pre-deployment, and enabling responsible disclosure in a manner that balances organizational and community interests. However, promising approaches, such as cooperative audits (Wilson et al. 2021), participatory audits (Dennler et al. 2023) and AI red teaming (OpenAI 2023; Ganguli et al. 2022), are beginning to take shape as the stakes surrounding ML systems continue to grow. This section traces the disclosure of ethical issues in machine learning, covering its history and trends.
Reporting has primarily relied on affected communities facing issues and then raising complaints through social media campaigns, media coverage, lawsuits, and other public pressure tactics after the fact. There have been few systematic mechanisms for auditing and identifying model issues proactively during development phases or continuously monitoring systems post-deployment. Prominent examples include recidivism prediction algorithms like COMPAS exhibiting significant racial bias against Black defendants, only uncovered through investigative journalism efforts (ProPublica 2016). Facial analysis systems misclassifying people of color often go unnoticed until public backlash erupts on social platforms (BBC News 2015). Harmful biases in language models generating toxic outputs against marginalized groups are frequently observed only due to vigilance within directly impacted communities (Cosmos Magazine 2023; Abid, Farooqi, and Zou 2021).
The lack of established processes enabling coordinated discovery and disclosure of ethical issues in machine learning and the closed-source nature of many AI systems means users are unaware how long problems fester before coming to light. The reliance on reactive reporting by harmed parties has made accountability an exception rather than the norm for ML systems. Researchers and practitioners alike have highlighted a pressing need to complement one-off investigations with proactive, community-centered mechanisms for responsible disclosures. A key way to enable decentralized discovery and disclosure of ML ethical flaws include providing legal protections for independent ML issue researchers, in a similar fashion to existing legal protections for good faith security research (Hacking Policy Council 2023).

### v3

# Coordinated Flaw Disclosure for AI: Beyond Security Vulnerabilities

@@ Abstract
Harm reporting in Artificial Intelligence (AI) currently lacks a structured process for disclosing and addressing algorithmic flaws, relying largely on an ad-hoc approach. This contrasts sharply with the well-established Coordinated Vulnerability Disclosure (CVD) ecosystem in software security. While global efforts to establish frameworks for AI transparency and collaboration are underway, the unique challenges presented by machine learning (ML) models demand a specialized approach. To address this gap, we propose implementing a Coordinated Flaw Disclosure (CFD) framework tailored to the complexities of ML and AI issues. This paper reviews the evolution of ML disclosure practices, from ad hoc reporting to emerging participatory auditing methods, and compares them with cybersecurity norms. Our framework introduces innovations such as extended model cards, dynamic scope expansion, an independent adjudication panel, and an automated verification process. We also outline a forthcoming real-world pilot of CFD. We argue that CFD could significantly enhance public trust in AI systems. By balancing organizational and community interests, CFD aims to improve AI accountability in a rapidly evolving technological landscape.
@@ 1 Introduction
The landscape of AI accountability is rapidly evolving, as exemplified by The New York Times’ lawsuit against OpenAI and Microsoft on December 27, 2023. This legal action alleges that ChatGPT was trained on the Times’ published works without authorization, providing evidence of verbatim article generation (Mac 2023). As part of a growing trend of litigation against AI providers (O’Brien 2023; Authors Guild 2023; Fenwick & West LLP 2023; ACLU 2022), this case highlights the complexities of replicating and verifying such claims. The Times’ lawsuit has garnered significant attention, partly due to its financial resources and media reach - advantages often unavailable to smaller affected parties. This underscores how such factors can shape public discourse and potentially drive improvements in AI systems through external pressure (Raji and Buolamwini 2022).
The emergence of unexpected model behaviors beyond intended scope and purpose emphasizes the need for robust frameworks to identify and address these issues effectively. This stands in contrast to the well-established Common Vulnerabilities and Exposures (CVE) ecosystem in cybersecurity, where software vulnerabilities are systematically cataloged and addressed with greater clarity and consensus.
@@
Historically, AI accountability initiatives have relied heavily on periodic audits, emphasizing repetitive assessments but lacking a structured framework for reporting user-identified issues post-deployment. Recent policy efforts have begun to acknowledge this gap (European Commission 2020). This audit-centric paradigm is reflected in influential policies such as the U.S. Executive Order on AI (The White House 2023), the EU AI Act (European Parliament and Council 2024), and New York City’s Local Law 144 (New York City Department of Consumer and Worker Protection 2021). However, this approach falls short when compared to the more comprehensive Coordinated Vulnerability Disclosure (CVD) processes standard in software security.
@@
CVD plays a crucial role in enabling independent researchers to report newly identified vulnerabilities to affected vendors and the public (MultiMedia LLC 1999). This process facilitates transparent remediation before potential exploitation by malicious actors and has become a vital practice enshrined in government regulations and industry standards. Notably, the FDA mandates the implementation of CVD programs for medical device companies to enhance cybersecurity (U.S. Food and Drug Administration 2023).
@@
While CVD has proven effective in traditional software security, its direct application to machine learning (ML) systems faces unique challenges. ML diverges from traditional software in two key aspects: 1) ML model issues must satisfy statistical validity thresholds, and 2) concerns related to trustworthiness and bias often extend beyond the typical scope of security vulnerabilities (Office of Science and Technology Policy 2022). Many critical bias reports discussed in social media and news may not align with the accepted definition of a vulnerability [^1] The Coordinated Vulnerabilities Enumeration program, managed by MITRE and central to the CVD ecosystem, defines a vulnerability as: A problem in software, firmware, hardware, or service component resulting from a weakness that can be exploited, causing a negative impact on the confidentiality, integrity, or availability of the affected component or components..
@@
Recognizing this limitation in the conventional definition of a vulnerability, we propose expanding the terminology with a new term - Flaw. We define a flaw as any unexpected model behavior that is outside of the defined intent and scope of the model design.
@@
To address the unique challenges posed by ML systems, we advocate for adapting the CVD process into a dedicated “Coordinated Flaw Disclosure” (CFD) framework. Tailored to ML’s distinctive properties, CFD aims to formalize the recognition of valid issues in ML models through an adjudication process. This process seeks to balance the interests of vendors and the broader community, enhancing public trust by curating disclosed flaws into actionable reports. The CFD framework introduces several key innovations:
  • Extended model cards that include detailed documentation of intent and scope, providing a baseline for flaw assessment.
  • A dynamic scope expansion mechanism that allows for the inclusion of common, unforeseen uses of the model.
  • An independent adjudication panel to mediate disputes and handle edge cases.
  • An automated verification process to streamline issue reporting and reproduction.
@@
In the subsequent sections of this paper, we explore specific aspects of this proposal. Section 2 provides essential background information, comparing periodic audits and ad hoc harm reporting while exploring emerging participatory auditing practices. Section 3 examines evolving trends in disclosure practices within ML systems and identifies their shortcomings. Section 4 outlines the proposed changes within the CFD framework to improve disclosure rigor. Section 5 explores the practical implementation of CFD, proposing the utilization of model cards to define scope and intent. We conclude by discussing limitations of our framework and future work which includes a real-world implementation pilot and the development of a Common Use Enumeration (CUE) system to standardize the tracking of AI model applications.
@@ 2 Background
In the following, we describe the background of our work. We will use the terminology of the glossary in Table 1.
@@
Intent The stated purpose, goals, and intended use cases of an AI model or system, as defined by its creators or deployers. This includes the model’s primary functions, target audience, and expected performance metrics.
Scope The boundaries of an AI model’s intended operation, including what the model is designed to do, what it is not designed to do, and known limitations or exclusions. It defines the range of inputs, outputs, and use cases that are considered valid for the model.
@@ 2.1 Coordinated Disclosure Conventions in Cybersecurity
CVD operates within a broader framework of programs that identify, index, and communicate security vulnerabilities, including the National Vulnerability Database, CVE system, and Common Vulnerability Scoring System (CVSS). These programs support diverse security operations such as vulnerability management, risk assessments, penetration testing, patch management, and threat intelligence (Cybersecurity Coalition 2019). CVE Numbering Authorities (CNA) are organizations external to MITRE that are authorized to issue CVEs within their scope, but MITRE acts as a “root” that handles disputes. For example vulnerabilities in open source projects managed by the Apache Software Foundation will be sent to them first and then, if needed, appealed to MITRE (Apache 2024). In the EU, the European Cybersecurity Scheme on Common Criteria (EUCC) is currently developed to cover information and communications technology (ICT) based on the international Common Criteria scheme (European Union Agency for Cybersecurity (2024) ENISA).
@@ 3 Disclosure Practices in Machine Learning Systems: Evolving Trends
This section examines trends in disclosure practices for machine learning (ML) systems and the work needed to harmonize these conventions. We analyze two axes: objective security and privacy flaw disclosures, and subjective ethical flaw disclosures, such as societal bias.
@@ 3.1 Security Disclosures for ML Systems
ML systems face three main types of disclosures: security vulnerability reports, model design failures, and error reports.
@@
These reports highlight security vulnerabilities in ML models, often in components around the model. Examples include CVE reports on Numpy (CVE 2019a), Tensorflow (CVE 2019b), and Pytorch (CVE 2023c). The OpenAI bug bounty on Bugcrowd addresses infrastructure flaws (OpenAI 2023). SafeTensors by Hugging Face (Hugging Face 2022) is an emerging effort to address reproducibility concerns.
An instance of a traditional vulnerability is DeepSloth (Hong et al. 2020), a Denial Of Service attack on LLMs. ProofPudding (MITRE 2019) was the first CVE issued for an issue caused by an ML model. The mitigation involved limiting API information (Oksuz, Halimi, and Ayday 2023).
@@
Model design failures manipulate samples to induce changes in output. Unlike infrastructure reports, they rely on modifying model output. Examples include the “Bayesian Poisoning Attack” (The MIT 2004 Spam Conference ) and the Cylance malware detector (Cyber 2019). The Malware Detection Evasion Competition (Anderson 2019) highlights such vulnerabilities.
Adversarial examples are often considered low priority (Szegedy et al. 2014). They can be defeated by simple techniques like jpeg compression (Aydemir, Temizel, and Temizel 2018) or blur & repair (Carlini et al. 2022a). While adversarial patches (Brown et al. 2018) are possible, they are hard to execute (Athalye et al. 2018). Marines bypassed an ML model with a simple attack (Scharre 2023b), highlighting that lower-hanging fruit is often more appealing (Apruzzese et al. 2022).
@@
Error reports stem from expected model errors or systematic errors amendable through data changes. These include:
  • One-off Errors: Normal errors within the model’s expected error rate.
  • Systematic Errors: Errors from dataset biases, leading to harmful outputs. Examples include the Google image classifier’s mislabeling (Scharre 2023a) and the ChatGPT divergence attack (Nasr et al. 2023).
Reports that Depend on Context
Recent reports include prompt injection attacks (NVIDIA 2023), with vulnerabilities in LangChain receiving CVEs. These can be “model design failure reports” or “error reports” depending on the context. For instance, basic chatbots might not be vulnerable, but complex chatbots must have measures to address these risks (Burgess 2023). Proper disclosure is crucial for systems granting sensitive access, such as email or calendar functions.
@@ 3.2 Disclosure of Ethical Issues in ML Systems
Despite calls for audits from various stakeholders (Raji et al. 2022; Fjeld et al. 2020; Wilkinson et al. 2023), the identification and disclosure of ethical issues in machine learning (ML) systems have been ad hoc and reactive. Unlike traditional software systems with established practices for coordinated vulnerability disclosure, ML model issues often surface inconsistently. Without a mandated reporting mechanism, harms caused by flawed ML systems are often discovered post-damage (Scharre 2023a). Few proactive mechanisms exist for auditing these complex systems, identifying pre-deployment risks, and enabling responsible disclosure balancing organizational and community interests. However, approaches like cooperative audits (Wilson et al. 2021), participatory audits (Dennler et al. 2023), and AI red teaming (OpenAI 2023; Ganguli et al. 2022) are emerging as ML stakes grow. This section traces the history and trends in disclosing ethical issues in ML.
Reporting has primarily relied on affected communities raising complaints through social media, media coverage, lawsuits, and public pressure after issues arise. Few systematic mechanisms audit and identify model issues proactively. Examples include the racial bias in COMPAS recidivism prediction algorithms uncovered by investigative journalism (ProPublica 2016) and facial analysis systems misclassifying people of color until public backlash erupts (BBC News 2015). Harmful biases in language models generating toxic outputs against marginalized groups are often observed due to vigilance within directly impacted communities (Cosmos Magazine 2023; Abid, Farooqi, and Zou 2021).
The lack of coordinated discovery and disclosure processes and the closed-source nature of many AI systems mean users remain unaware of problems until they surface. Reactive reporting by harmed parties has made accountability an exception rather than the norm for ML systems. Researchers and practitioners emphasize the need for proactive, community-centered mechanisms for responsible disclosures. Providing legal protections for independent ML issue researchers, akin to protections for good faith security research, is essential (Hacking Policy Council 2023).